If Wild Widgets, Inc., were an all-equity company, it would have a beta of 1.20. The company has a target debt-equity ratio of .70. The expected return on the market portfolio is 10 percent and Treasury bills currently yield 3.5 percent. The company has one bond issue outstanding that matures in 30 years, a par value of $1,000, and a coupon rate of 6.4 percent. The bond currently sells for $1,070. The corporate tax rate is 22 percent.

  

a.

What is the company’s cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the company’s cost of equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What is the company’s weighted average cost of capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a

Cost of debt
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =30
1070 =∑ [(6.4*1000/100)/(1 + YTM/100)^k]     +   1000/(1 + YTM/100)^30
                   k=1
YTM = 5.90

b

Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (expected return on the market - risk-free rate)
Cost of equity% = 3.5 + 1.2 * (10 - 3.5)
Cost of equity% = 11.3

c

D/A = D/(E+D)
D/A = 0.7/(1+0.7)
=0.4118
Weight of equity = 1-D/A
Weight of equity = 1-0.4118
W(E)=0.5882
Weight of debt = D/A
Weight of debt = 0.4118
W(D)=0.4118
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 5.8970471427*(1-0.22)
= 4.599696771306
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=4.6*0.4118+11.3*0.5882
WACC =8.54%